Thirty-five small communities in Connecticut (population near 10,000 each) gave an average of x = 138.5 reported cases of larceny per year. Assume that σ is known to be 44.5 cases per year.
(a) Find a 90% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
lower limit
Incorrect: Your answer is incorrect.
upper limit
Incorrect: Your answer is incorrect.
margin of error
Incorrect: Your answer is incorrect.
(b) Find a 95% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
lower limit
upper limit
margin of error
(c) Find a 99% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
lower limit
upper limit
margin of error

Answer by stanbon(75887)

Thirty-five small communities in Connecticut (population near 10,000 each) gave an average of x = 138.5 reported cases of larceny per year. Assume that σ is known to be 44.5 cases per year.
(a) Find a 90% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
x-bar = 138.5
ME = 1.645*44.5/sqrt(35) = 12.38
-----------------------------
lower limit = 138.5-12.38 = 126.1
-----------------
upper limit = 138.5+12.38 = 150.0
------------------------------------------

(b) Find a 95% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
lower limit:: 138.5-1.96*44.5/sqrt(35) = 123.8
upper limit:: 138.5+1.96*44.5/sqrt(35) = 153.2
============================
(c) Find a 99% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
Comment:: Change the "z" value to 2.5758 and do the arithmetic
--------------------
lower limit
upper limit
=====
margin of error
Cheers,
Stan H.
----------------